OpenStudy (anonymous):

2x-y=8 7x-3y=28 what is the solution..SO confused and frustrated...please help

OpenStudy (anonymous):

let's use elimination method of solving system of equations. This means we need the -y in the first equation to become +3y. So we must first multiply the first equation completely by -3

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

once we do that we have to add the equations together.

OpenStudy (anonymous):

but its asking me to use the substitution method and express as an ordered pair

OpenStudy (anonymous):

isolate the y in the first equation

OpenStudy (anonymous):

the ordered pair will be found once we know x, and y. (x,y)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2x-y=8 -y=8-2x y=-8+2x substitute y for -8+2x in the second equation to solve for x

OpenStudy (anonymous):

7x-3y=28 7x-3(-8+2x)=28 you can work out the rest by yourself yes?